OpenStudy (anonymous):

A square field has an area of 479ft^2. What is the approximate length of a side of the field? Give your answer to the nearest foot. A. 21 B. 240 C.23 D. 22

Directrix (directrix):

The area of a square is the square of one of its congruent sides. A = s^2 479 = s^2 21.88 = s To the nearest foot, 22 = s